# A Comparative Study on Positional Encoding for Time-frequency Domain Dual-path Transformer-based Source Separation Models

### v1

@@ IV-A Datasets
WHAMR! dataset [26] contains monaural noisy reverberant two-speaker mixtures sampled at 8 kHz. Speech and noise signals are sourced from WSJ0 [27] and WHAM! [28], respectively. The dataset comprises 20000, 5000, and 3000 mixtures for training, validation, and testing, respectively. The model is trained to perform dereverberation, denoising, and separation simultaneously.
MUSDB18-HQ contains 100 and 50 stereo songs sampled at 44.1 kHz for training and testing, respectively. For data partitioning, we adopted a commonly used split, allocating 86 songs for training and 14 for validation. In MUSDB, each song has four stems (vocals, bass, drums, and other), and the goal is to separate the mixture into its individual stems.
@@ IV-D Results on WHAMR!
TABLE I: SI-SDRi on WHAMR! dataset. ‘min‘ version is used for training.
@@
TABLE II: SI-SDRi when $K=1$ on WHAMR! dataset. Training is done with an 8kHz ‘min‘ version, while testing is done using both ‘min‘ and ‘max‘ versions. Training chunk was 4-seconds.
@@
Table I presents the SI-SDR improvement (SI-SDRi) achieved by TF-Locoformer trained with different PEs. Both A* and B* used the 8kHz min version of training data; however, A* models were trained with one-second input segments, whereas B* models were trained with four-second input segments.
As shown in Table I, the NoPE approach achieves performance comparable to or better than other explicit PE methods under matched conditions (8k min). In contrast, Table II shows the results for models without convolution, where the convolutional kernel size $K$ is set to one, demonstrating that explicit PEs outperform NoPE on 8k min data. These results suggest that in models incorporating convolutional layers, such as TF-Locoformer, convolution implicitly encodes position information, allowing explicit PE to be omitted without compromising overall performance. Notably, in our experimental setup, training B4 (NoPE) was approximately 16% faster than training B3 (RoPE). To further assess whether NoPE maintains its effectiveness in larger models, we evaluated a medium-sized TF-Locoformer, with results summarized in Table III. The NoPE-based model outperformed the RoPE-based model and other state-of-the-art models, demonstrating its scalability.

### v2

@@ IV-A Datasets
WHAMR! [26] contains noisy reverberant two-speaker mixtures sampled at 8 kHz. Only the first channel was used to evaluate performance in a monaural setup. Speech and noise signals are sourced from WSJ0 [27] and WHAM! [28], respectively. The dataset comprises 20000, 5000, and 3000 mixtures for training, validation, and testing, respectively. The model is trained to perform dereverberation, denoising, and separation simultaneously.
MUSDB18-HQ contains 100 and 50 stereo songs sampled at 44.1 kHz for training and testing, respectively. For data partitioning, we adopted a commonly used split, allocating 86 songs for training and 14 for validation. The goal is to separate the mixture into the individual stems (vocals, bass, drums, and other). The model takes stereo mixtures as input and estimates stereo-separated signals.
@@ IV-D Results on WHAMR!
TABLE I: Average SI-SDRi on WHAMR! test set. ‘min‘ version is used for training, while evaluation is done on both ‘min‘ and ‘max‘ versions, each with 8kHz and 16kHz data.
@@
TABLE II: Average SI-SDRi when $K=1$ on WHAMR! test set. Training is done with 8kHz ‘min‘ version, while testing is done using both ‘min‘ and ‘max‘ versions. Training chunk was 4-seconds.
@@
Table I presents the average SI-SDR improvement (SI-SDRi) on the WHAMR! test set. Both A* and B* used the 8kHz min version of training data; however, A* models were trained with one-second input segments, whereas B* models were trained with four-second input segments.
As shown in Table I, the NoPE approach achieves performance comparable to or better than other explicit PE methods under matched conditions (8k min). In contrast, Table II shows the results for models without convolution, where the convolutional kernel size $K$ is set to one, demonstrating that some explicit PEs outperform NoPE on 8k min data. These results suggest that in models incorporating convolutional layers, such as TF-Locoformer, convolution implicitly encodes position information, allowing explicit PE to be omitted without compromising overall performance. Notably, in our experimental setup, training B4 (NoPE) was approximately 16% faster than training B3 (RoPE). To further assess whether NoPE maintains its effectiveness in larger models, we evaluated a medium-sized TF-Locoformer, with results summarized in Table III. The NoPE-based model outperformed the RoPE-based model and other state-of-the-art models, demonstrating its scalability.
@@ V Conclusion
VI Acknowledgement
This work was supported by JSPS KAKENHI Grant Number JP24KJ2096.